OpenStudy (anonymous):

2.25(4x-4)=-2+10x+12

OpenStudy (anonymous):

2.25(4x-4) = 9x - 9 So 9x - 9 = -2 + 10x + 12 9x - 9 = 10x + 10

OpenStudy (anonymous):

its it idetity and does it have solution or no solution

OpenStudy (anonymous):

It has solution. Subtract both side by 9x